OpenStudy (anonymous):

prove that 2/y is equal to 1/y^2 using ln

OpenStudy (anonymous):

it isnt

OpenStudy (whpalmer4):

Well, it usually isn't, but it is at \[\frac{2}{y}=\frac{1}{y^2}\]\[2y=1\]\[y=\frac{1}{2}\]

OpenStudy (anonymous):

I think he typed up an incomplete question as the "ln" seems to be as if he was in the middle of typing Induction. @melmel